A financial analyst is in the process of reviewing four investments projects for one of his clients. The net present cash values for the four projects are estimated at $50 million, $15 million, $20 million and $80 million respectively. His client is committed to $20 million of capital funds for the first year and $13 million for the second year. The cash requirements for each project in the two-year period are as follows:

       Projects -->

Period i

v

1

2

3

4

First year

6

2

4

10

Second year

8

4

8

6

Net present value

$50 (million)

$15(million)

$20(million)

$80(million)

Formulate for the financial analyst.Describe the optimal solution to the client.
Answer #1

All projects having positive NPVs, thus all projects are feasible.

(All figures are in $' million)

Funds required to invest in all projects are

First year = 6 + 2 + 4 + 10 = 22 & available fund for first year is only 20.

Second year = 8 + 4 + 8 + 6 = 36 & available fund for second year is only 13.

In these type of situations we use Profitability Index to decide which projects are selected and which are to be skipped.

Profitablilty index = PV of cash inflow/ PV of cash outflows

But in this such information is not given to calculate Profitability index, thus we are calculating here NPV per One $ of investment.

thus NPV per One $ of investment = NPV of project / Investment in Project

Note: We are taking here value of investment in project for both two year with out taking effect of time value of money as no discount rate is provided in the question.

Project NPV

Investment in two year

NPV per one $ of invsetment Priority
1 50 14 3.57 II
2 15 6 2.5 III
3 20 12 1.67 IV
4 80 16 5 I

Total fund available with investor = 20+13 = 33

Total fund required for Project 4 & Project 1= 16 + 14 =30

thus he can invest in only project 4 & Project 1, for investing in next profitable project i.e. project 2 he requires $6 million but he has only $3 million in his hands.

Thus the optimal solution for the client is to invest in Project 4 & Project 1.

Thus Funds available in first year = 20, Investment in First year = 10+6 = 16, Funds remains in hand =4

Funds available in second year = 4+ 13= 17, Investment in second year =6+8= 14, funds remains in hand = 3

NPV from total investment = 80 + 50 = 130